OpenStudy (pythagoras123):

There are 20 questions in a quiz. 5 marks are awarded for each correct answer, 2 marks are deducted for each wrong answer and 1 mark was awarded if the question was not attempted. Sophia got 67 marks altogether. How many questions did she get correct?

OpenStudy (anonymous):

assumeing that 15 correct ans...

OpenStudy (anonymous):

she got 14 right 3 wrong and skipped 3

OpenStudy (anonymous):

this makes total score of 67 :D

OpenStudy (anonymous):

good logic @hashsam1

OpenStudy (anonymous):

14*5=70 right answers 3*2=6 wrong ones 70-6=64 as she had skipped 3 thus she was awarded 3 marks her total score 64+3=67 Thanks

OpenStudy (marco26):

Hmm, I guess this is the equation ,, :) not sure, though 5x + y -2z = 67 x + y + z = 20 Let x be the no. of correct answer, y be the skiipped answer, z be the wrong no...and solve,,:)

OpenStudy (anonymous):

@marco26 can you solve that for us? :P

OpenStudy (marco26):

this is consistent with hashsam1 's answer

OpenStudy (marco26):

lol,, this can be done by gauss jordan reduction, or any convenient method,..anyway, the answer is already there :)